OpenStudy (anonymous):

. How many solutions are there to the following system of equations? 4x - 14y = 6 -2x + 7y = -3 A. 0 B.1 C.2 D.infinitely many

OpenStudy (cwrw238):

hint: multiply the second equation by -2 what do you get?

OpenStudy (therealmeeeee):

ok -2(4x^2+4x+1) a=4; b=4; c=1

OpenStudy (cwrw238):

? is this answer from another question?

OpenStudy (therealmeeeee):

so its B

OpenStudy (therealmeeeee):

no dude

OpenStudy (cwrw238):

?? - I'm going crazy here is the question about 2 equations 4x - 14y = 6 and -2x + 7y = -3?

OpenStudy (cwrw238):

??

OpenStudy (therealmeeeee):

its B @Miz.BeautiNation

OpenStudy (cwrw238):

1?

OpenStudy (anonymous):

4x - 14y = 6 (change to slope intercept form ) -14y = -4x + 6 y = -4x/-14 + 6/-14 y = 2/7 - 3/7 -2x + 7y = -3 (change to slope intercept form) 7y = 2x - 3 y = 2/7x - 3/7 since both equations are the same in slope intercept form, then there are INFINITE SOLUTIONS

OpenStudy (anonymous):

oops...I forgot to put the x in the first equation.....y = 2/7x - 3/7